Given that p(x) =k÷x2 is a probability distribution for a random variable that can take on the values x = 0, 1, 2, 3, and 4.

a. Find k.

b. Find the expression for the cdf F(x) of the random variable. 



1
Expert's answer
2021-07-04T17:29:19-0400

a. Finding K

"\sum_{i=0}^4P(Xi)=1"

The PMF should be "\frac{k}{2^x}" and not "\frac{k}{x^2}" since we cannot divide a number by zero.

"\sum_{i=0}^4P(Xi)=k+\frac{k}{2}+\frac{k}{4}+\frac{k}{8}+\frac{k}{16}=\frac{31k}{16}"

"\frac{31k}{16}=1" Solving for k,

"k=\frac{16}{31}"

b. CDF

"CDF=\sum_0^xP(x)" Substituting the value of k in "\sum_{i=0}^4P(Xi)" above, we get;

"\sum_0^4P(x)=\frac{16}{31}+\frac{8}{31}+\frac{4}{31}+\frac{2}{31}+\frac{1}{31}"

Note that it forms a geometric progression with "a=\frac{16}{31}" and "r=\frac{1}{2}". Note also that it consist of 5 terms since the value of x starts from 0.

The sum of n terms of a GP when "r<1" is

"S_n=\frac{a(1-r^n)}{1-r}"

But to allow 0 to 4 values of x, n becomes x+1.

"CDF=\frac{\frac{16}{31}(1-\frac{1}{2^{x+1}})}{1-\frac{1}{2}}" Simplifying, we have

"CDF=\frac{32}{31}(1-\frac{1}{2^{x+1}})"